# Generalized mKdV Equation and Genus Two Jacobi Type Hyperelliptic Differential Equation

### v1

@@
We generalized the mKdV equation in order that the static equations include $\rm sn$} differential equation. As a result, a good correspondence was obtained between the KdV equation and the mKdV equation. For general genus two hyperelliptic curves, we obtained differential equations for Weierstrass type and Jacobi type hyperelliptic functions. Considering the special case of $\lambda_{6}=0,\lambda_{0}=0$, Weierstrass type and Jacobi type hyperelliptic functions are different solutions to the same hyperelliptic differential equations. Then these solutions are connected by the special $\rm Sp(4,{\bf R})$} Lie group transformation.
@@ 1 Introduction
The reason why soliton system can be solved exactly despite being nonlinear is that there is the Lie group structure. It is well known that the KdV equation has the ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [1, 2]. We have demonstrated that $\wp$ function, which is the special static solution of the KdV equation, has the ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [3], and the genus two hyperelliptic $\wp$ functions have the ${\rm Sp(4,{\bf R})/Z_{2}}\cong{\rm SO(3,2)}$ Lie group structure [3, 4, 5]. Not in the general Riemann surfaces but in the hyperelliptic curves, we expect that multi-periodic functions, i.e., hyperelliptic functions have the Lie group structure.
In this paper, we first generalized mKdV equation in order that the static differential equation includes $\rm sn$} differential equation. This gives a good correspondence between the KdV equation and the mKdV equation. For general genus two hyperelliptic curves, we obtain the conventional differential equations for Weierstrass type and Jacobi type hyperelliptic functions. If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for Weierstrass type and Jacobi type hyperelliptic functions are different solutions of the same hyperelliptic differential equations. We confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.
@@ 2.1 Integrability condition for the generalized mKdV equation
where we denote $\dfrac{\partial u(x,t)}{\partial t}=u_{t}(x,t)$, $\dfrac{\partial u(x,t)}{\partial x}=u_{x}(x,t)$, $\dfrac{\partial^{3}u(x,t)}{\partial x^{3}}=u_{xxx}(x,t)\rule$[-9.95845pt]{0.0pt}{25.60747pt} etc. Eq.(2.2) is generalized in order that the static equation includes differential equation for the $\rm sn$} function as follows
@@
$\eta$ is the spectral parameter and $b$ is a constant parameter, which is determined later. The matrix $M$ in Eq.(2.21) has $\rm GL(2,{\bf R})$} structure. The integrability condition of Eqs.(2.1) and (2.2) is given by
@@
and $b$ is set to be $a=4\eta^{2}(b-1)$. Then the above integrability condition $D=0$ gives the generalized mKdV equation.
@@ 2.2 Various transformations between the static KdV and the static generalized mKdV equations
For simplicity, we will denote $v(x,0)$ as $v(x)$. This type of abbreviation is used in other cases as well.
@@ 2.2.1 Miura transformation
Of course, the time dependent KdV and generalized mKdV equations are linked by the time dependent Miura transformation $u(x,t)=v^{2}(x,t)+v_{x}(x,t)-a/6$ [8].
@@ 2.2.3 Inverse square transformation
In this case, $v(x)$ not only satisfies the static generalized mKdV equation but also becomes the Jacobi $\rm sn$} function. This transformation is well-known that links the Weierstrass $\wp$ function and the Jacobi $\rm sn$} function in the form
@@ 2.2.4 Inverse power transformation
In this case, $v(x)$ is a solution of the static generalized mKdV equation, but it is not that of the Jacobi $\rm sn$} function.
@@ 2.2.5 A relationship between transformations
where $v_{1}(x)$ and $v_{2}(x)$ satisfies the same $\rm sn$} type differential equation. In order that Eq.(2.40) becomes the standard $\rm sn$} differential equation, we put $k=\sqrt{2}$, $a=(1+k^{2})/2=3/2$, and we obtain [3]
@@
On the other hand, ${\widehat{{\rm sn}}(z)=\frac{1}{k\,{\rm sn}(z)}}$ satisfies the same differential equation
@@ 3 Differential equations for genus two Jacobi type hyperelliptic functions
Though $\Re_{22}$ and $\Re_{21}$ are symmetric functions, but they are not the fundamental symmetric functions. Then any symmetric functions cannot be expressed in the closed form by $\Re_{22}$ and $\Re_{21}$. So far, closed differential equations for the genus two hyperelliptic functions are known only for the $\lambda_{6}=0$ case [10].
@@ 3.1 Genus two Weierstrass type hyperelliptic differential equations
We call the above as the Weierstrass type hyperelliptic functions. Here $\wp_{22}$ and $\wp_{21}$ are fundamental symmetric function. Then any symmetric quantities can be expressed by $\wp_{22}$ and $\wp_{21}$. Thus we can obtain the hyperelliptic differential equations in the closed forms. Though we obtain the relation
@@
the condition ${\frac{\partial\wp_{22}}{\partial u_{1}}=\frac{\partial\wp_{21}}{\partial u_{2}}}$ alone is sufficient to guarantee the integrability condition, i.e. the existence of $\sigma(u_{1},u_{2})$ function as a potential. If $\wp_{22}$ and $\wp_{21}$ are given as functions of $\sigma$ function, $Q$ can in principle be given as a function of $\sigma$ function via the generalized Kummer surface relation. We will give the generalized Kummer surface relation later.
In order to obtain the differential equations for genus two Weierstrass type hyperelliptic functions, we first calculate $\wp_{2222}$, $\wp_{2221}$, $\wp_{2211}$, $\wp_{2111}$, $Q_{11}$, $Q_{21}$ and $Q_{22}$. Because these quantities are symmetric functions of $x_{1}$ and $x_{2}$, these are expressed by $\wp_{22}$ and $\wp_{21}$. If these quantities contain symmetric function of the type $y_{1}y_{2}$, the quantity $y_{1}y_{2}$ is expressed by $Q$ and/or $Q^{2}$ by using ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$,
@@
From Eq.(3.12), we naturally obtain $Q$ in the combination of Eq.(3.9).
The characteristic feature of $\lambda_{6}\neq 0$ hyperelliptic differential equations is the existence of the cubic term, that is, terms ${-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{3},-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}\wp_{21},\cdots,-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}Q}$, in Eq.(3.11) $\sim$ Eq.(3.17).
Here we give the generalized Kummer surface relation. For $\wp_{22}$, $\wp_{21}$ and $Q$, only two are independent. Then there exist one relation, which is called the generalized Kummer surface relation. By using the relation ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$, we obtain
@@
Using Eq.(3.1), this constraint is expressed by $\wp_{22}$, $\wp_{21}$ and $Q$. The explicit expression is given as follows. We denote $X=\wp_{22}$, $Y=\wp_{21}$, $Z=Q$, and we obtain the following generalized Kummer surface relation of the form
@@
If we consider the special case of $\lambda_{6}=0$, we obtain ${\frac{\partial\wp_{21}}{\partial u_{1}}=\frac{\partial Q}{\partial u_{2}}}$, then we can put $Q=\wp_{11}$. Thus we obtain the conventional differential equations for Weierstrass type hyperelliptic functions [11]. If we consider the more special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for Weierstrass type hyperelliptic differential functions become
@@
The condition $\lambda_{0}=0$ is not the special restriction, because we can always put $\lambda_{0}=0$ by the constant shift of $x_{i}$ in the hyperelliptic curve.
@@ 3.2 Genus two Jacobi type hyperelliptic differential equations
Next, we try to obtain differential equations for the genus two Jacobi type hyperelliptic functions by the following dual transformation. Under the dual transformation
@@
From Eq.(3.37), we naturally obtain $\widehat{Q}$ in the combination of Eq.(3.33). The characteristic feature of the $\lambda_{0}\neq 0$ hyperelliptic differential equations is the existence of the cubic terms ${-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{3},-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{\wp}_{11},\cdots,-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{Q}}$, in Eq.(3.30) $\sim$ Eq.(3.35). .
If we take the special case of $\lambda_{6}=0$, differential equations for the genus two Weierstrass type hyperelliptic functions does not contain cubic terms just as the differential equation of the genus one Weierstrass $\wp$ function. While the above differential equations for genus two Jacobi type hyperelliptic functions contain cubic terms just as the differential equation of the genus one Jacobi $\rm sn$} function. This is the reason why we call $\widehat{\wp}_{ij}$, $\widehat{Q}$ as the Jacobi type hyperelliptic functions.
In the more special case of $\lambda_{0}=0,\lambda_{6}=0$, we obtain ${\frac{\partial\widehat{Q}}{\partial u_{1}}=\frac{\partial\widehat{\wp}_{21}}{\partial u_{2}}}$, then we can put $\widehat{Q}=\widehat{\wp}_{22}$ in such special case. Then the correspondence of Eq.(3.33) becomes $\wp_{22}\rightarrow\widehat{\wp}_{11},\wp_{21}\rightarrow\widehat{\wp}_{21},\wp_{11}\rightarrow\widehat{\wp}_{22}$. In this special case, differential equations for the genus two Jacobi type hyperelliptic equations become
@@
in genus one case. Eq.(3.47) is the special case of the half-period relation of the standard projective type II $\rm Sp(4,{\bf R})$} Lie group transformation
@@ 4 Summary and Discussions
The static KdV equation has solutions of the Weierstrass $\wp$ function. The Weierstrass $\wp$ elliptic functions and the Jacobi $\rm sn$} elliptic functions form a family of elliptic functions. We have generalized the mKdV equation in such a way as the static mKdV differential equation and the static KdV differential equation have good correspondence. We have demonstrated that this generalized mKdV equation is an integrable equation. We have also demonstrated that there are four kinds of relations between solutions of the static KdV equation and the generalized static mKdV equation. Next, we generalized the relationship between $\wp$ and $\rm sn$} functions to the genus two case.
For the general genus two hyperelliptic curves, we obtained differential equations for the genus two Weierstrass type and Jacobi type hyperelliptic functions.
If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, the differential equations of the genus two Weierstrass type and Jacobi type hyperelliptic functions become the same. Then the Weierstrass type and the Jacobi type hyperelliptic functions are linked as the half-period shift solutions of the same differential equations. In fact, we have confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.

### v2

@@
We generalized the mKdV equation such that the static equations included the $\rm sn$} differential equation. As a result, a good correspondence was obtained between the KdV and mKdV equations. For the general genus two hyperelliptic curves, we obtained differential equations for the Weierstrass-type and Jacobi-type hyperelliptic functions. Considering the special case of $\lambda_{6}=0,\lambda_{0}=0$, the Weierstrass-type and Jacobi-type hyperelliptic functions are different solutions to the same hyperelliptic differential equations. These solutions are then connected via a special $\rm Sp(4,{\bf R})$} Lie group transformation.
@@ 1 Introduction
The soliton system can be solved exactly despite being nonlinear because of the Lie group structure. It is well known that the KdV equation has an ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [1, 2]. We have demonstrated that $\wp$ function, which is the special static solution of the KdV equation, has the ${\rm Sp(2,{\bf R})/Z_{2}}\cong{\rm SO(2,1)}$ Lie group structure [3], and the genus two hyperelliptic $\wp$ functions have the ${\rm Sp(4,{\bf R})/Z_{2}}\cong{\rm SO(3,2)}$ Lie group structure [3, 4, 5]. In the hyperelliptic curves, not in general Riemann surfaces, we expect that multi-periodic functions, that is, hyperelliptic functions, have a Lie group structure.
In this study, we first generalized the mKdV equation such that the static differential equation included the $\rm sn$} differential equation. This gives a good correspondence between the KdV and mKdV equations. For the general genus two hyperelliptic curves, we obtain the conventional differential equations for the Weierstrass-type and Jacobi-type hyperelliptic functions. If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, differential equations for the Weierstrass-type and Jacobi-type hyperelliptic functions are different solutions of the same hyperelliptic differential equations. We confirmed that the two hyperelliptic functions were linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.
@@ 2.1 Integrability condition for the generalized mKdV equation
where we denote $\dfrac{\partial u(x,t)}{\partial t}=u_{t}(x,t)$, $\dfrac{\partial u(x,t)}{\partial x}=u_{x}(x,t)$, $\dfrac{\partial^{3}u(x,t)}{\partial x^{3}}=u_{xxx}(x,t)\rule$[-9.95845pt]{0.0pt}{25.60747pt} etc. Eq.(2.2) is generalized such that the static equation includes a differential equation for the $\rm sn$} function as follows
@@
$\eta$ is the spectral parameter and $b$ is a constant parameter, which is determined later. The matrix $M$ in Eq.(2.21) has $\rm GL(2,{\bf R})$} structure. The integrability condition for Eq.(2.12) and (2.21) are given by
@@
and $b$ is set to be $a=4\eta^{2}(b-1)$. Then the above integrability condition $D=0$ yields the generalized mKdV equation.
@@ 2.2 Various transformations between the static KdV and the static generalized mKdV equations
For simplicity, we will denote $v(x,0)$ as $v(x)$. This type of abbreviation is also used in other cases.
@@ 2.2.1 Miura transformation
The time-dependent KdV and generalized mKdV equations are linked by the time-dependent Miura transformation $u(x,t)=v^{2}(x,t)+v_{x}(x,t)-a/6$ [8].
@@ 2.2.3 Inverse square transformation
In this case, $v(x)$ not only satisfies the static generalized mKdV equation, but also becomes the Jacobi $\rm sn$} function. This transformation is well-known to link the Weierstrass $\wp$ function and Jacobi $\rm sn$} function in the form
@@ 2.2.4 Inverse power transformation
In this case, $v(x)$ is a solution of the static generalized mKdV equation, but not that of the Jacobi $\rm sn$} function.
@@ 2.2.5 A relationship between transformations
where $v_{1}(x)$ and $v_{2}(x)$ satisfy the same $\rm sn$}-type differential equation. In order that Eq.(2.40) becomes the standard $\rm sn$} differential equation, we set $k=\sqrt{2}$, $a=(1+k^{2})/2=3/2$, and obtain
@@
On the other hand, ${\widehat{{\rm sn}}(z)=\frac{1}{k\,{\rm sn}(z)}}$ satisfies the same differential equation [3]
@@ 3 Differential equations for genus two Jacobi type hyperelliptic functions
Though $\Re_{22}$ and $\Re_{21}$ are symmetric functions, but they are not the fundamental symmetric functions. Therefore, it is not always possible to express any symmetric function in a closed form using $\Re_{22}$ and $\Re_{21}$. Until now, closed differential equations for the genus two hyperelliptic functions are known only for the $\lambda_{6}=0$ case [10].
@@ 3.1 Genus two Weierstrass type hyperelliptic differential equations
We refer to this as Weierstrass-type hyperelliptic functions. Here $\wp_{22}$ and $\wp_{21}$ are fundamental symmetric function. Then any symmetric quantity can be expressed by $\wp_{22}$ and $\wp_{21}$. Thus, we can obtain hyperelliptic differential equations in closed forms. Though we obtain the relation
@@
the condition ${\frac{\partial\wp_{22}}{\partial u_{1}}=\frac{\partial\wp_{21}}{\partial u_{2}}}$ alone is sufficient to guarantee the integrability condition, that is, the existence of $\sigma(u_{1},u_{2})$ function as a potential. If $\wp_{22}$ and $\wp_{21}$ are given as functions of $\sigma$ function, $Q$ can in principle be given as a function of $\sigma$ function via the generalized Kummer surface relation. We will provide a generalized Kummer surface relation later.
In order to obtain the differential equations for genus two Weierstrass type hyperelliptic functions, we first calculate $\wp_{2222}$, $\wp_{2221}$, $\wp_{2211}$, $\wp_{2111}$, $Q_{11}$, $Q_{21}$ and $Q_{22}$. Because these quantities are symmetric functions of $x_{1}$ and $x_{2}$, these are expressed by $\wp_{22}$ and $\wp_{21}$. If these quantities contain symmetric function of the type $y_{1}y_{2}$, the quantity $y_{1}y_{2}$ is expressed by $Q$ and/or $Q^{2}$ by using ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$.
@@
From Eq.(3.12), $Q$ naturally appears as a combination of the terms in Eq.(3.9). The characteristic feature of $\lambda_{6}\neq 0$ hyperelliptic differential equations is the existence of the cubic term, that is, terms ${-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{3},-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}\wp_{21},\cdots,-\frac{32\lambda_{6}}{\lambda_{5}^{2}}\wp_{22}^{2}Q}$, in Eq.(3.11) $\sim$ Eq.(3.17).
Here we present a generalized Kummer surface relation. For $\wp_{22}$, $\wp_{21}$ and $Q$, only two are independent. Then, there exists one relation, called the generalized Kummer surface relation. Using the relation ${y_{1}y_{2}=-2(x_{1}-x_{2})^{2}Q+\frac{1}{2}F(x_{1},x_{2})}$, we obtain
@@
Using Eq.(3.1), this constraint is expressed by $\wp_{22}$, $\wp_{21}$ and $Q$. The explicit expression is as follows. We denote $X=\wp_{22}$, $Y=\wp_{21}$, $Z=Q$, and obtain the following generalized Kummer surface relation of the form
@@
If we consider the special case of $\lambda_{6}=0$, we obtain ${\frac{\partial\wp_{21}}{\partial u_{1}}=\frac{\partial Q}{\partial u_{2}}}$, then we can put $Q=\wp_{11}$. Thus, we obtain conventional differential equations for Weierstrass-type hyperelliptic functions [11]. If we consider the more special case of $\lambda_{6}=0,\lambda_{0}=0$, the differential equations for the Weierstrass-type hyperelliptic differential functions become
@@
The condition $\lambda_{0}=0$ is not a special restriction, because we can always put $\lambda_{0}=0$ by the constant shift of $x_{i}$ in the hyperelliptic curve.
@@ 3.2 Genus two Jacobi type hyperelliptic differential equations
Next, we try to obtain differential equations for the genus two Jacobi-type hyperelliptic functions using the following dual transformation. Under a dual transformation
@@
From Eq.(3.37), $\widehat{Q}$ naturally appears as a combination of the terms in Eq.(3.33). The characteristic feature of the $\lambda_{0}\neq 0$ hyperelliptic differential equations is the existence of the cubic terms ${-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{3},-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{\wp}_{11},\cdots,-\frac{32\lambda_{0}}{\lambda_{1}^{2}}\widehat{\wp}_{21}^{2}\widehat{Q}}$, in Eq.(3.35) $\sim$ Eq.(3.41).
If we take the special case of $\lambda_{6}=0$, differential equations for the genus two Weierstrass type hyperelliptic functions does not contain cubic terms just as the differential equation of the genus one Weierstrass $\wp$ function. The above differential equations for genus two Jacobi-type hyperelliptic functions contain cubic terms just as the differential equation of the genus one Jacobi $\rm sn$} function. This is why we call $\widehat{\wp}_{ij}$, $\widehat{Q}$ Jacobi-type hyperelliptic functions.
In the special case of $\lambda_{0}=0,\lambda_{6}=0$, we obtain ${\frac{\partial\widehat{Q}}{\partial u_{1}}=\frac{\partial\widehat{\wp}_{21}}{\partial u_{2}}}$, then we can put $\widehat{Q}=\widehat{\wp}_{22}$ in such a special case. Therefore, the correspondence in Eq.(3.33) is as follows: $\wp_{22}\rightarrow\widehat{\wp}_{11},\wp_{21}\rightarrow\widehat{\wp}_{21},\wp_{11}\rightarrow\widehat{\wp}_{22}$. In this special case, differential equations for the genus two Jacobi type hyperelliptic equations become
@@
in genus one case. Eq.(3.47) is a special case of the half-period relation of the standard projective type II $\rm Sp(4,{\bf R})$} Lie group transformation
@@ 4 Summary and Discussions
The static KdV equation has solutions of the Weierstrass $\wp$ function. The Weierstrass $\wp$ elliptic function and Jacobi $\rm sn$} elliptic functions form a family of elliptic functions. We have generalized the mKdV equation in such a way that the static mKdV differential equation and static KdV differential equation have good correspondence. We demonstrated that this generalized mKdV equation is integrable. We have also demonstrated that there are four kinds of relations between the solutions of the static KdV equation and the generalized static mKdV equation. Next, we generalized the relationship between $\wp$ and $\rm sn$} functions to the genus two case.
For the general genus two hyperelliptic curves, we obtained differential equations for the genus two Weierstrass-type and Jacobi-type hyperelliptic functions.
If we consider the special case of $\lambda_{6}=0,\lambda_{0}=0$, the differential equations of the genus two Weierstrass-type and Jacobi-type hyperelliptic functions become the same. Then the Weierstrass-type and the Jacobi-type hyperelliptic functions are linked as half-period shift solutions of the same differential equations. In fact, we confirmed that two hyperelliptic functions are linked by a special half-period $\rm Sp(4,{\bf R})$} Lie group transformation.